OpenStudy (anonymous):

In the equation x – 4y = 2, for what value of x will y = 0

geerky42 (geerky42):

plug 0 in y, then solve for x.

geerky42 (geerky42):

x - 4y = 2 x - 4(0) = 2 ... ?

OpenStudy (anonymous):

6 thank you

geerky42 (geerky42):

None. what is 4 times 0?

OpenStudy (anonymous):

0

geerky42 (geerky42):

Right, so x - 4(0) = 2 -> x - 0 = 2 What is value of x?

OpenStudy (anonymous):

2

geerky42 (geerky42):

Yep

geerky42 (geerky42):

That's your answer.